(x-7)/(square root x-1) + 3(x-1)^1/2=0
What does 'x' equal?

Answer by ad_alta(240)

If you mean to solve (x-7)/(x-1)^{1/2} + 3(x-1)^{1/2}=0, the easiest way to proceed is to multiply both sides by (x-1)^{1/2}. That gives 4x-10=0, so x=5/2. If you mean (x-7)/(x^{1/2}-1) + 3(x-1)^{1/2}=0 the best way is to multiply by the square root conjugate (x^{1/2}+1), although the equation is still quite nasty. I think you probably meant the former.